OpenStudy (df001):

Sin(x)-csc(x)/csc(x) = -cos^2(x)

OpenStudy (df001):

Help how do i solve this?

OpenStudy (df001):

hi please help me

OpenStudy (df001):

\[\frac{ \sin x - \csc x}{ \csc x } = - \cos ^{2} x\]

OpenStudy (alekos):

whats another way of expressing cosecx ?

OpenStudy (df001):

1/sinx?

OpenStudy (alekos):

so replace cosecx with 1/sinx

OpenStudy (df001):